OpenStudy (anonymous):

what is the formula for v given, KE=1/2mv^2?

OpenStudy (tanya123):

v = (2*100J/10kg)^1/2 = (20 m^2s^-2)^1/2 = 4.47 m/s

OpenStudy (anonymous):

without given amount of KE and m

OpenStudy (anonymous):

@tanya123

OpenStudy (tanya123):

To start, we simply rearrange the equation so we have velocity as a function of mass and energy. When we do this we get v = (2KE/m)^1/2 Now we plug in the values to obtain a solution v = (2*100J/10kg)^1/2 = (20 m^2s^-2)^1/2 = 4.47 m/s